OpenStudy (anonymous):

Solve x2 + 6x + 7 = 0

OpenStudy (anonymous):

\[-b \pm \frac{ \sqrt{b^2-4ac} }{2a }\] Apply that, remember your equation is ax^2+bx+c=0

OpenStudy (anonymous):

x = –1 and x = –5 ??

OpenStudy (anonymous):

?

OpenStudy (jiteshmeghwal9):

here a=1, b=6 & c=7. \[\Huge{-(6) \pm \sqrt{(6)^2-4(1)(7)}\over2(1)}\]

OpenStudy (anonymous):